# PooDLe[Uncaptioned image]: Pooled and dense self-supervised learning from naturalistic videos

### v1

@@ 4 Experiments
We pretrain PooDLe on raw videos from BDD100K [50] and Walking Tours (WT) [42] and evaluate them with semantic segmentation and object detection benchmarks. The BDD100K pretrained model is evaluated on the corresponding in-domain benchmarks as well as Cityscapes [14] and the Walking Tours model on ADE20K [1]. We also ablate our combination of loss functions and decoder components as well as the effects of crop area and input resolutions.
@@ 4.1 Experiment Setup
1) BDD [50] consists of 100,000 dashcam driving videos collected in various weather conditions and times of day from New York and the San Francisco Bay Area. Each video is  40 seconds long at 720p and 30 fps. We pretrain with the 70,000 videos in the official training split and evaluate on the semantic segmentation and object detection tasks proposed alongside the dataset. 2) Walking Tours (WT) [42] is a dataset of first-person egocentric videos collected from YouTube of a continuous walkaround through various cities of Europe, Asia, and a wildlife safari. There are 10 videos, ranging from 59 minutes to 2 hours 55 minutes, at 720p and 30 fps. Each video contains large numbers of unique objects per frame and natural transitions in lighting and location. For self-supervised pretraining, we use either the Venice video ($\text{WT}_{\text{Venice}}$) or all 10 videos ($\text{WT}_{\text{all}}$) following DoRA [42].
@@
We use ResNet-50 (R50) [23] as our feature encoder, with the dense projector and predictor networks following FlowE [49] and pooled counterparts following BYOL [20]. For SDM, we use two decoder stages, with each consisting of a 2×2$\times$2 × upsample, a ResNet Bottleneck Block [23] for computation, and a 2-layer convolutional MLP for the lateral connection. When training on BDD, we sample two frames that are $0.5\sim 1$ seconds apart ($\Delta t\in\{15...30\}$) from each video. We then take two large crops from the same image coordinates of area $[0.16,0.45]$ of the original image and resize them to $512\times 1024$ pixels before applying augmentations. For each training epoch on WT, we divide each video into 10-second clips and randomly sample two frames $0.5$ seconds apart from each clip, and use crop area range $[0.65,1.0]$. For non-spatial augmentations on both datasets, we apply color distortion and Gaussian blurring independently to each frame following BYOL [20]. For the dense objective, we also apply a random reversible affine transformations similar to FlowE [49]: random scaling of 0.9–1.1$\times$ and rotation of -10–10. For the local objective, we sample $K=6$ subcrop pairs and use a crop area of $[0.05,0.3]$ of the initial dense crop for both BDD and WT, resized to $192\times 192$. When sampling supcrops, random spatial jitter is ±$\pm 10\%$ of the large crops’ height and width. We attempt experiments with ViT-S, but observe a collapse phenomenon (see Appendix G).
@@
We use official implementations of DenseCL, PixPro, DINO, iBOT and DoRA, and our own implementation of FlowE for pretraining on BDD. We use torchvision for ImageNet supervised and weights released online for ImageNet-pretrained DINO. We obtain weights from the authors of DoRA for iBOT, DINO - ViT, and DoRA pretrained on WT and use official implementations of DINO - R50 and PixPro for pretraining on WT. For PixPro, we use its FPN decoder only in pretraining, which we find is best for performance. We use 512x1024 crops for training FlowE and DINO-R50; we are unable to train the other baselines with high-resolution crops due to prohibitive GPU memory requirements.
@@
We adopt the evaluations described in FlowE [49] for BDD and Cityscapes. We use DeepLab v1 [10] as the ‘linear’ readout header and UperNet [47] as the heavier readout head for semantic segmentation and Faster R-CNN with ResNet-C4 and Faster R-CNN with FPN [24] as the standard and heavier readout headers for object detection. We do not include ViT object detection due to the lack of an existing reliable recipe. For semantic segmentation on ADE20K, we perform both linear readout following BDD and Upernet finetuning as described in iBOT [54]. We retain the SDM when evaluating PooDLe on semantic segmentation with linear readouts. We report mean intersection-over-union (mIoU), pixel-level accuracy (Acc), and mean average precision (mAP) as our evaluation metrics. Additional details on implementation and hyperparameters are provided in Appendix A.
@@ 4.2 Main Results
Table 1: BDD semantic segmentation (SemSeg) and object detection (Det) readout evaluations. All settings are conducted with a frozen backbone. *Pretrained on BDD, initialized with supervised ImageNet weights.
BDD100K Sem. Seg. BDD100K Obj. Det. Cityscapes Sem. Seg
@@
DenseCL [45] R50 800 BDD 21.3 82.6 46.9 91.4 0.9 23.4 25.1 84.8 57.2 93.0
FlowE [49] R50 100 BDD 35.7 88.5 47.3 91.5 3.2 23.8 43.1 89.5 57.7 93.1
PixPro [48] R50 100 BDD 21.8 80.0 37.3 88.0 0.7 18.4 25.5 81.0 44.3 89.5
DINO [7] R50 100 BDD 13.1 64.7 25.6 80.3 0.3 11.9 14.9 69.4 29.2 81.4
DINO [7] ViT-S 300 BDD 29.6 86.8 41.1 90.1 - - 35.1 87.9 51.5 91.9
iBOT [54] ViT-S 800 BDD 27.2 85.4 35.5 88.7 - - 32.0 86.2 44.0 90.3
DoRA [42] ViT-S 200 BDD 33.2 88.1 43.3 90.7 - - 37.4 88.7 50.8 92.0
@@
We evaluate the learned representations on semantic-segmentation and object-detection on the BDD in-distribution, labeled benchmark (Table 1). PooDLe achieves superior performance on all readout tasks compared to other models pretrained on BDD, outperforming the strongest baseline FlowE by $3.5\%$ mIoU on linear and $2.6\%$ mIoU on UperNet for semantic segmentation, and $1.7\%$ mAP on C4 and $1.4\%$ mAP on FPN for object detection. This indicates that PooDLe is able to learn better representations for dense prediction tasks. While we keep the decoder in linear readout, it is discarded during UperNet probing, showing that the representation improves substantially over all other BDD-trained models ($49.9$ mIoU compared to next-best $47.3$). We demonstrate in Table 3 that our performance improvements are derived from better recognition of small classes. We also evaluate the downstream transfer of PooDLe representations by performing linear and UperNet readout on the Cityscapes benchmark with models trained on BDD. PooDLe readily outperforms all baselines. We provide visualizations of predicted segmentation masks in Figure 4 and additional evaluation visualizations in Figure 12 and Figure 13.
@@
PooDLe outperforms ImageNet supervised pretraining as well, despite the latter’s advantage in learning small and rare classes present in BDD100K (spatial imbalance shown in Figure 1) due to ImageNet being a class-balanced dataset with iconic views of objects. In addition, we may pretrain PooDLe on BDD100K with weights initialized from the ImageNet supervised checkpoint to further improve performance, notably boosting mIoU by $8\%$ and Acc by $6\%$ on linear semantic segmentation. We provide additional comparisons to ImageNet-pretrained baselines in Appendix E, where PooDLe remains competitive despite being pretrained in the challenging naturalistic video setting.
@@
Table 2: ADE20K semantic segmentation linear readout and finetuning evaluations. Linear readout is performed with a frozen backbone while in finetuning, backbone parameters are trainable. DINO-ViT and iBOT results are taken from DoRA [42].
Method Arch Epoch Pretrain SemSeg Linear Finetune
mIoU Acc mIoU Acc
DINO [7] R50 800 IN1K 15.7 61.5 43.0 80.5
DINO [7] ViT-S 100 IN1K - - 33.9 -
PixPro [48] R50 100 $\text{WT}_{\text{Venice}}$ 4.6 48.6 36.0 77.6
DINO [7] R50 100 $\text{WT}_{\text{Venice}}$ 6.9 48.2 35.7 77.4
DINO [7] ViT-S 100 $\text{WT}_{\text{Venice}}$ 7.8 57.7 29.2 74.7
iBOT [54] ViT-S 100 $\text{WT}_{\text{Venice}}$ 33.9 -
DoRA [42] ViT-S 100 $\text{WT}_{\text{Venice}}$ 14.1 63.5 35.2 77.7
PooDLe R50 20 $\text{WT}_{\text{Venice}}$ 14.6 59.0 36.6 77.9
DoRA [42] ViT-S 100 $\text{WT}_{\text{all}}$ 13.9 64.4 38.3 79.3
PooDLe R50 20 $\text{WT}_{\text{all}}$ 16.5 63.9 41.0 79.6
WT-pretrained models.
We also train PooDLe on $\text{WT}_{\text{Venice}}$ as well as the entire dataset, $\text{WT}_{\text{all}}$. Table 2 shows linear readout and finetuning experiments on ADE20K [53] semantic segmentation following Venkataramanan et al. [42]. Notably, when pretrained on $\text{WT}_{\text{all}}$, PooDLe surpasses DoRA by $2.6\%$ mIoU on linear readout and $2.7\%$ mIoU on UperNet finetuning over DoRA. PooDLe also shows some improvement trained on $\text{WT}_{\text{Venice}}$-pretraining setting, with a gain of $0.5\%$ mIoU on linear readout over DoRA and $0.6\%$ mIoU on UperNet finetuning over PixPro. This shows that PooDLe also learns strong representations from naturalistic video captured in the open world. We show visualizations of predicted segmentation masks for ADE20K evaluations in Figure 14.
@@
Naturalistic videos have severe class and size imbalance (Figure 1; e.g., for semantic segmentation, the "road" occupies $21\%$ of pixels and appears in $96\%$ of images while "bicycle" only occupies $0.05\%$ in $6\%$ of images. Capturing information on severely underrepresented classes is very challenging. While it is unfeasible to balance our training data, we demonstrated the composability of  PooDLe with iconic pretraining by initializing with supervised ImageNet weights (Table 1). This variant improves upon the semantic class understanding of IN1K-pretraining while also significantly improving on spatial boundaries (accuracy), which we hypothesize is due to our dense loss $\mathcal{L}_{\text{dense}}$. In particular, training PooDLe on BDD improves performance on linear readout by $8\%$ mIoU and $6\%$ Acc over the supervised IN1K weights used for initialization.
To further evidence this phenomenon, we designate BDD classes as "small" if they on average occupy $<1\%$ of pixels and "large" as those that occupy $>1\%$. We also independently define "rare" as classes that appear in $<20\%$ of images and "common" as those that appear $>20\%$. Table 3 shows linear readout mIoU with different class groupings to further illustrate the class and spatial imbalance effect. Full class-level statistics and designations are in Appendix F. We observe that FlowE performs well on large and common classes due to its dense loss, but struggles on small or rare classes. Meanwhile, the supervised ImageNet model effectively learns about smaller classes via it’s balanced pretraining data as evidenced by its performance on semantic segmentation. PooDLe, with its unified objectives, and spatial decoder module shows significant improvement over other BDD-trained models in all class designations and particularly on small and rare classes. We also train a PooDLe on top of supervised ImageNet weights. It drastically improves on large classes, from $42.2\%$ to $56.1\%$, due to the dense objective, while remaining competitive with supervised ImageNet on small classes.j
@@ 4.3 Ablation studies
Table 4.2 shows our ablation experiments, testing each of our contributions beginning from FlowE. Models trained without the decoder use dilated convolutions in place of pooling operations, as in FlowE [49]. How the dense and pooled objectives are composed with and without the decoder is shown in Figure 3. For this table, models are trained for $40$ epochs on BDD and use a reduced $256\times 512$ resolution and $[0.04,0.11]$ area for the initial crops; we evaluate on BDD semantic segmentation using linear readout.
@@
(a) Global crop area
(b) Subcrop area
(c) Resolution
Figure 5: Experiments varying image cropping parameters for global and subcrops. (a) vary global crop area (b) varying subcrop area with effect on large and small class groups (c) jointly varying global corp area and input resolution to keep constant pixel density.
@@
We also demonstrate that PooDLe is able to perform well even with self-supervised flow. We train our own unsupervised UFlow [30] model on KITTI [18] and finetune it on BDD resulting in only a 0.5% mIoU loss compared to pretraining with RAFT. See Appendix H and A for visualizations and details.

### v2

@@ 4 Experiments
We pretrain PooDLe on raw videos from BDD100K (Yu et al., 2020) and Walking Tours (WT) (Venkataramanan et al., 2024) and evaluate them on semantic segmentation and object detection benchmarks. The BDD100K pretrained model is evaluated on in-distribution tasks as well as Cityscapes (Cordts et al., 2016), and the Walking Tours model on ADE20K (Agrawal et al., 2015) and our newly proposed Walking Tours Semantic benchmark. We also ablate our combination of loss functions and decoder components, as well as the effects of crop area and input resolutions.
@@ 4.1 Experiment Setup
1) BDD (Yu et al., 2020) consists of 100,000 dashcam driving videos collected in various weather conditions and times of day from New York and the San Francisco Bay Area. Each video is ~40 seconds long at 720p and 30 fps. We pretrain with the 70,000 videos in the training split and evaluate on the dataset’s semantic segmentation and object detection tasks. 2) Walking Tours (WT) (Venkataramanan et al., 2024) is a dataset of first-person YouTube videos of a continuous walkaround through various cities of Europe, Asia, and a wildlife safari. There are 10 videos, ranging from 59 minutes to 2 hours 55 minutes, at 720p and 30 fps. Each video contains large numbers of unique objects per frame and natural transitions in lighting and location. We use either the Venice video ($\text{WT}_{\text{Venice}}$) or all 10 videos ($\text{WT}_{\text{all}}$) following DoRA (Venkataramanan et al., 2024).
@@
We use ResNet-50 (R50) (He et al., 2016a) as our feature encoder, with the dense projector and predictor networks following FlowE (Xiong et al., 2021) and pooled counterparts following BYOL (Grill et al., 2020). For SDM, we use two decoder stages, with each consisting of a 2×2$\times$2 × upsample, a ResNet Bottleneck Block (He et al., 2016a), and a 2-layer convolutional MLP for the lateral connection. When training on BDD, we sample two frames that are $0.5\sim 1$ seconds apart ($\Delta t\in\{15...30\}$) from each video. We then take two large crops from the same image coordinates of area $[0.16,0.45]$ of the original image and resize them to $512\times 1024$ pixels before applying augmentations. For each training epoch on WT, we divide each video into 10-second clips and randomly sample two frames $0.5$ seconds apart from each clip, and use crop area range $[0.65,1.0]$. For both datasets, we apply color distortion and Gaussian blurring independently to each frame following BYOL (Grill et al., 2020). For the dense objective, we also apply random reversible affine transformations similar to FlowE (Xiong et al., 2021): random scaling of 0.9–1.1$\times$ and rotation of -10–10 degrees. For the local objective, we sample $K=6$ subcrop pairs with a crop area of $[0.05,0.3]$ of the initial crop, resized to $192\times 192$ for both BDD and WT. For subcrops, random spatial jitter is applied as ±$\pm 10\%$ of the initial crops’ height and width.
@@
We use official implementations of DenseCL, PixPro, DINO, iBOT, and DoRA, and our own implementation of FlowE for pretraining on BDD. We use torchvision for supervised ImageNet (IN1K) and weights released online for ImageNet-pretrained DINO. We obtain weights from the authors of DoRA for iBOT, DINO-ViT, and DoRA pretrained on WT and use official implementations of DINO-R50, MAE, and PixPro for pretraining on WT. For PixPro, we use either its FPN decoder or high-resolution crops for pretraining and report results from the best-performing setting. We use $512\times 1024$ crops to train all R50 baselines for more accurate comparisons.
@@
We adopt the evaluation protocol from FlowE (Xiong et al., 2021) for BDD and Cityscapes. We use DeepLab v1 (Chen et al., 2018) as the “linear” readout header and UperNet (Xiao et al., 2018) as the heavier readout head for semantic segmentation and Faster R-CNN with ResNet-C4 and Faster R-CNN with FPN (He et al., 2016b) as the standard and heavier readout headers for object detection. We do not include ViT object detection due to the lack of an established recipe. For semantic segmentation on ADE20K, we perform both linear readout following BDD and UperNet finetuning as described in iBOT (Zhou et al., 2021). We retain the SDM when evaluating PooDLe on semantic segmentation with linear readouts but discard it when using UperNet. We report mean intersection-over-union (mIoU), pixel-level accuracy (Acc), and mean average precision (mAP) as our evaluation metrics. Additional details on implementation and hyperparameters are provided in Appendix A.
@@ 4.2 Main Results
Table 1: BDD and CityScapes semantic segmentation (SemSeg) and object detection (Det) readout evaluations. All settings are conducted with a frozen backbone. *Pretrained on BDD, initialized with supervised IN1K weights.
BDD100K Sem. Seg. BDD100K Obj. Det. Cityscapes Sem. Seg.
@@
DINO (Caron et al., 2021) ViT-S 300 BDD 29.6 86.8 41.1 90.1 - - 35.1 87.9 51.5 91.9
iBOT (Zhou et al., 2021) ViT-S 800 BDD 27.2 85.4 35.5 88.7 - - 32.0 86.2 44.0 90.3
DoRA (Venkataramanan et al., 2024) ViT-S 200 BDD 33.2 88.1 43.3 90.7 - - 37.4 88.7 50.8 92.0
DINO (Caron et al., 2021) R50 100 BDD 13.1 64.7 25.6 80.3 0.3 11.9 14.9 69.4 29.2 81.4
PixPro (Xie et al., 2021) R50 100 BDD 21.8 80.0 37.3 88.0 0.7 18.4 25.5 81.0 44.3 89.5
DenseCL (Wang et al., 2021b) R50 100 BDD 24.2 84.9 41.8 90.0 0.7 20.3 26.6 85.6 53.2 91.9
FlowE (Xiong et al., 2021) R50 100 BDD 35.7 88.5 47.3 91.5 3.2 23.8 43.1 89.5 57.7 93.1
@@
We report our results on semantic segmentation and object detection on the BDD100K benchmark in Table 1. PooDLe achieves superior performance on all readout tasks compared to prior methods, outperforming the strongest baseline FlowE by $3.5\%$ mIoU on linear and $2.6\%$ mIoU on UperNet for semantic segmentation, and $1.7\%$ mAP on C4 and $1.4\%$ mAP on FPN for object detection. We find that that PooDLe’s improved performance (Table 3) is attributed to better recognition of small and rare object classes. We also evaluate the transfer of PooDLe representations to new tasks by evaluating on the Cityscapes benchmark, where PooDLe outperforms all baselines. Figure 4 shows predicted segmentation masks, and Figure 14 and Figure 15 show additional evaluation results.
@@
PooDLe also outperforms supervised IN1K pretraining, despite the latter’s advantage in learning small and rare classes present in BDD100K (spatial imbalance shown in Figure 1) due to ImageNet being a class-balanced dataset with iconic views of objects. In addition, we pretrain PooDLe on BDD100K with weights initialized from the supervised IN1K checkpoint, improving mIoU by $8\%$ and Acc by $6\%$ over the initialization weights on linear semantic segmentation. In Appendix F, we show PooDLe remains competitive against IN1K-pretrained baselines despite being trained in the challenging naturalistic video setting.
@@
Table 2: ADE20K and WT-Sem semantic segmentation linear readout and finetuning evaluations. Linear readout is performed with a frozen backbone while in finetuning, backbone parameters are trainable. DINO-ViT and iBOT results are taken from DoRA (Venkataramanan et al., 2024).
ADE20K Sem. Seg. WT-Sem Sem. Seg.
Method Arch Epoch Pretrain SemSeg Linear Finetune SemSeg Linear Finetune
mIoU Acc mIoU Acc mIoU Acc mIoU Acc
DINO (Caron et al., 2021) R50 800 IN1K 15.7 61.5 43.0 80.5 8.8 76.7 17.8 87.6
DINO (Caron et al., 2021) ViT-S 100 IN1K - - 33.9 - - - - -
DINO (Caron et al., 2021) ViT-S 100 $\text{WT}_{\text{Venice}}$ 7.8 57.7 29.2 74.7 4.6 73.7 11.0 83.0
iBOT (Zhou et al., 2021) ViT-S 100 $\text{WT}_{\text{Venice}}$ 33.9 - - - - -
MAE (He et al., 2022) ViT-S 100 $\text{WT}_{\text{Venice}}$ 7.4 55.1 24.1 71.4 4.3 72.6 8.9 81.5
DoRA (Venkataramanan et al., 2024) ViT-S 100 $\text{WT}_{\text{Venice}}$ 14.1 63.5 35.2 77.7 6.2 76.9 13.6 85.7
DINO (Caron et al., 2021) R50 100 $\text{WT}_{\text{Venice}}$ 6.9 48.2 35.7 77.4 4.2 69.0 12.3 84.7
PixPro (Xie et al., 2021) R50 100 $\text{WT}_{\text{Venice}}$ 4.6 48.6 36.0 77.6 3.7 69.3 11.5 84.2
PooDLe R50 20 $\text{WT}_{\text{Venice}}$ 14.6 59.0 36.6 77.9 6.4 75.7 13.7 85.4
DINO (Caron et al., 2021) ViT-S 100 $\text{WT}_{\text{all}}$ 34.1 - - - - -
MAE (He et al., 2022) ViT-S 100 $\text{WT}_{\text{all}}$ 10.6 60.4 31.4 75.9 6.6 77.7 12.7 85.2
DoRA (Venkataramanan et al., 2024) ViT-S 100 $\text{WT}_{\text{all}}$ 13.9 64.4 38.3 79.3 7.8 79.4 15.9 87.5
PooDLe R50 20 $\text{WT}_{\text{all}}$ 16.5 63.9 41.0 79.6 11.2 81.3 17.0 86.9
@@
WT-pretrained models.
We also train PooDLe on $\text{WT}_{\text{Venice}}$ and $\text{WT}_{\text{all}}$. Table 2 shows results on ADE20K (Zhou et al., 2017) semantic segmentation using linear readout and finetuning following (Venkataramanan et al., 2024). Notably, when pretrained on $\text{WT}_{\text{all}}$, PooDLe obtains $2.6\%$ higher mIoU than DoRA on ADE20K linear readout, and $2.7\%$ mIoU on UperNet finetuning. PooDLe also performs better on $\text{WT}_{\text{Venice}}$, with a gain of $1.4\%$ mIoU over DoRA and $0.6\%$ mIoU over PixPro on ADE20K UperNet finetuning. Figure 16 shows predicted segmentation masks for ADE20K.
Walking Tours Semantic benchmark.
While ADE20K is a challenging benchmark, it contains a mixture of indoor and outdoor scenes that can be out-of-distribution from scenes in Walking Tours. Therefore, we introduce Walking Tours Semantic (WT-Sem) to provide a more in-distribution benchmark to accompany the WT dataset. We find that when pretrained on $\text{WT}_{\text{all}}$, PooDLe outperforms DoRA (Venkataramanan et al., 2024) by $3.4\%$ and $1.1\%$ mIoU on linear readout and UperNet finetuning, respectively. To generate the dataset, we use OpenSeeD (Zhang et al., 2023a), a strong open-vocabulary segmentation model, to generate semantic segmentation masks for all videos in $\text{WT}_{\text{all}}$ as well as 3 new walkaround videos. We use the Swin-L (Liu et al., 2021) variant of OpenSeeD finetuned on ADE20K semantic segmentation with a vocabulary of the 150 class labels from ADE20K to generate masks. See Appendix D for visualizations and details of WT-Sem.
@@
classes
Subcrop area small large all
0.04 - 0.18 16.0 41.6 34.6
0.18 - 0.36 14.2 41.2 33.7
0.36 - 0.54 13.6 39.5 32.4
0.54 - 1.00 12.3 38.8 31.5
Table 5: Choice of subcrop area on small, large and all classes.
Figure 5: Varying input resolution and global crop area as a fraction of the full frame. Large resolutions prefer larger crops.
@@
Naturalistic videos have imbalanced class representation and object sizes (Figure 1; e.g., “road” occupies $21\%$ of pixels while “bicycle” only occupies $0.05\%$ of pixels). Capturing information on these underrepresented classes is very challenging. To further demonstrate this phenomenon, we categorize BDD classes as “small” if they occupy $<1\%$ of pixels and “large” for those that occupy $>1\%$. Separately, we define “rare” as classes that appear in $<20\%$ of images and “common” as those that appear in $>20\%$. Table 3 shows linear readout mIoU for different class groupings, highlighting the impact of class and spatial imbalance. Full class-level statistics and designations are in Appendix G. We observe that FlowE performs well on large and common classes due to its dense loss, but struggles on small or rare classes. Meanwhile, supervised IN1K, benefiting from balanced pretraining data, effectively learns about smaller classes. PooDLe, with its unified objectives and spatial decoder module, significantly outperforms other BDD-pretrained models across all class groupings, particularly on small and rare classes. PooDLe, initialized from supervised IN1K weights, significantly improves upon supervised IN1K on large classes, from $42.2\%$ to $56.1\%$, due to the dense objective, while remaining competitive with supervised IN1K on small classes.
@@ 4.3 Ablation studies
Table 4.2 shows ablation experiments, testing each of our contributions beginning from FlowE. Models trained without the decoder use dilated convolutions in place of pooling operations, as in FlowE (Xiong et al., 2021). Figure 3 shows how the dense and pooled objectives are composed with and without the decoder. For the ablations, models are trained for $40$ epochs on BDD and use a reduced $256\times 512$ resolution and $[0.04,0.11]$ area for the initial crops; we evaluate on BDD semantic segmentation using linear readout.
@@
We also demonstrate that PooDLe is able to perform well even with self-supervised flow. We train UFlow (Jonschkowski et al., 2020) model on KITTI (Geiger et al., 2013) and finetune it on BDD resulting in only a 0.5% mIoU loss compared to pretraining with RAFT. See Appendix A and I for details and visualizations.